OpenStudy (anonymous):

How do you express the exact value of sin 75 as a signle fraction

terenzreignz (terenzreignz):

Use the simple fact that sin 75 = sin (30 + 45) Can you do it from here? :)

OpenStudy (anonymous):

Yeah, but I don't understand why I use 30 and 45. Is it because it's in the Chart (sin cos and tan) ?

terenzreignz (terenzreignz):

Yes, that's the reason :) They're the ones we have precise values for sin 30 = 1/2 sin 45 = sqrt(2)/2

terenzreignz (terenzreignz):

so knowing that, can you take over from here? :)

OpenStudy (anonymous):

yes, thanks!

terenzreignz (terenzreignz):

no problem :)